write an equation for the parabola with focus (0,0) and directrix y=6
please help

Answer by lwsshak3(11628)

write an equation for the parabola with focus (0,0) and directrix y=6
given data shows parabola opens downward:
Its basic form of equation: (x-h)^2=-4p(y-k), (h,k) =coordinates of the vertex
For given parabola:
axis of symmetry: x=0
x-coordinate of vertex=0
y-coordinate of vertex=3 (half the distance between focus and directrix on the axis of symmetry)
vertex:(0,3)
p=3 (distance from vertex to focus or directrix on the axis of symmetry
4p=12
equation of given parabola: x^2=-12(y-3)